What is the perimeter of a square if the diagonal is 24 ft?
Answer by richwmiller(17219)

a^2+b^2=c^2
where a=b and c^2=24
such that
2a^2=24
solve for a
this will give you one side of the square.
the perimeter of a square is 4a